# Not All Edges are Equally Robust: Evaluating the Robustness of Ranking-Based Federated Learning

### v1

@@ 1 Introduction
We evaluate our attack on five benchmark datasets and integrate eight Byzantine-robust aggregation rules (AGRs) within the FRL framework to test the attack impact of our VEM. The results demonstrate that our attack achieves a significantly greater attack impact compared to state-of-the-art attacks across all combinations of datasets, model architectures, and AGRs, with an overall attack impact of 53.23%, making it 3.7 times more effective than other attacks. Additionally, we developed a custom defense tailored specifically to counter our attack; however, our VEM attack remains effective against this targeted defense. Our key contributions can be summarized as follows:
  • We conduct the first systematic analysis of FRL’s robustness, uncovering a critical vulnerability within the framework. Our findings reveal that FRL does not guarantee the robustness of the system and there exists some edges that are particularly vulnerable to poisoning attack. Additionally, we provide a necessary condition to launch a successful attack in the FRL framework.
@@
  • We release the source code at https://anonymous.4open.science/r/VEM-E913 to facilitate future studies in this area.
@@ 2.3 Federated Rank Learning (FRL)
Lottery Tickets Hypothesis and Supermasks. Modern neural networks contain an enormous number of parameters, often leading to overparameterization [33, 34]. To solve this issue, Frankle and Carbin [16] offer an intriguing hypothesis i.e., Lottery Tickets Hypothesis (LTH), which states that within a randomly initialized, dense neural network (supernetwork), there exists a sparse subnetwork such that when trained in isolation, it can achieve the test accuracy as the fully trained original supernetwork within the same number of training iterations. Follow-up work by Ramanujan et al. [35] proposes the edge-popup (EP) algorithm [35] to effectively find subnetworks. Specifically, the EP algorithm assigns different scores to each edge in the supernetwork and uses these scores to determine which edges to retain. The process involves both a forward and a backward pass. During the forward pass, the EP selects the top kof edges with the highest scores. In the backward pass, it updates these scores using the straight-through gradient estimator [36], enabling gradients to pass through non-differentiable operations. The process is conducted iteratively to update the scores and retain the edges that effectively reduce the training loss. Detailed steps can be seen in Algorithm 2 in Appendix A.1. By doing so, it identifies a subnetwork that performs as well as the fully trained supernetwork.
@@ 3.2 Motivation
However, we observe that FRL is not inherently robust, and there are specific edges that are particularly vulnerable to poisoning attacks. To verify this argument, we record the reputation difference between each edge and the selection boundary $\gamma$, which is calculated as $\Delta_{j}=|w_{j}-\gamma|$, and $w_{j}$ is the aggregated reputation for edge $e_{j}$ (for $j\in[1,n]$). This metric indicates the minimal modification needed for edge $e_{j}$ to cross the selection boundary, thereby reflecting its vulnerability. A lower value of $\Delta_{j}$ signifies a higher vulnerability. As illustrated in Fig. 3, there is a significant variation in $\Delta_{j}$ (for $j\in[1,n]$), ranging from 0 to 5616 for the MNIST dataset and from 0 to 17208 for the CIFAR10 dataset. Additionally, we use the red dash line to indicate the max adversary damage (e.g., 20% malicious rate), the edges to the left of it are considered vulnerable edges.
@@ 3.3 VEM Overview
Our VEM aims to manipulate malicious updates such that the final global model deviates significantly from their original ones, thus negatively affecting the global model performance. To effectively achieve this goal and make the FRL framework tackleable, we first provide a formalization of FRL in Section 3.4. Specifically, we expand the 1D ranking into a 2D matrix to reduce the non-differentiable operations involved in the original algorithm.
@@ 3.5 Vulnerable Edges Identification
In this section, we provide a theoretical analysis of the range of vulnerable edges. Intuitively, if the reputation difference $\Delta_{j}$ between edge $e_{j}$ and the selection boundary is smaller than the maximum damage the adversary can cause, we call the edge $e_{j}$ a vulnerable edge. Formally, we have the following theorem.
@@
Implications. Given that the user input is discrete integers ranging from 1 to $n$, the adversary’s potential damage is limited by a maximum budget i.e., $m(a_{n}-a_{1})$ for $m$ malicious users. Therefore to launch a successful attack (i.e., move edges across the selection boundary), a necessary condition is that the needed reputation change, i.e., $\overline{w}_{\text{min}}-\overline{w}_{v}$ or $\overline{w}_{\text{max}}-\overline{w}_{v}$ should not exceed the adversary’s damage budget. Therefore, the edges that the adversary can alter (i.e., vulnerable edges) are bounded by Eq. (5). Moreover, as $m$ increases, the range of vulnerable edges expands. As shown in Fig. 6, the x-axis denotes the malicious rate and the y-axis indicates the percentage of vulnerable edges calculated based on Theorem 1. As the malicious rate rises, so does the percentage of vulnerable edges. When the malicious rate reaches 0.4, all edges can be considered vulnerable. Even with an extreme scenario where just 1 malicious client (4% malicious rate), there are still 11% and 10% vulnerable edges on the MNIST and CIFAR10 datasets, respectively.
@@
These two estimations are reasonable since in each round of the FRL training, clients get the global ranking from the previous round and train it only for a few local epochs (e.g., 2). Therefore, the $R_{u}(t)$ is still relatively similar to $R_{g}(t-1)$. Additionally, we conduct experiments to evaluate the performance of these estimation methods in benign cases. Specifically, we calculate the estimated $\overline{W}$ using these two methods and get the vulnerable edges set, namely $\mathbb{V}_{\text{alt}}$ and $\mathbb{V}_{\text{hist}}$. We then compare these sets with the true set of vulnerable edges $\mathbb{V}_{\text{true}}$ which is determined using full knowledge. Then we use estimation accuracy, which is calculated as calculated as $|\mathbb{V}_{\text{hist}}\cap\mathbb{V}_{\text{true}}|/|\mathbb{V}_{\text{true}}|$ and $|\mathbb{V}_{\text{alt}}\cap\mathbb{V}_{\text{true}}|/|\mathbb{V}_{\text{true}}|$ to measure the estimation performance. A higher value indicates a more accurate estimation. As illustrated in Fig. 7, where the y-axis represents the estimation accuracy and the x-axis depicts training iteration, both methods achieve a estimation accuracy of around 99.5% for the MNIST dataset and 99% for the CIFR10 dataset. Additionally, historical prediction demonstrates a slightly higher accuracy, which leads us to use it as our estimation method. Please refer to Appendix C to see more estimation performance under the non-IID setting.
Figure 7: Estimation accuracy during the FRL training.
@@ 3.6 Optimization
After identifying the vulnerable edges, we aim to target those edges and manipulate their ranking, thus effectively affecting the global model performance. Therefore, we have the following optimization function such that the global models reputation of those vulnerable edges deviate significantly from their original values. Formally, we have
@@ 4.1 Experimental setup
Datasets. We use four benchmark datasets widely used in prior works to evaluate our attacks, including MNIST [41], CIFAR10 [42], FashionMNIST [43], and EMNIST [44]. Details are deferred to Appendix B.1. For our main experiment, we use independently and identically distributed dataset (IID) because poisoning FL with IID data is the hardest [10, 11].
@@
Evaluation Metric. We follow [17] to use the attack impact to measure the attack effect. Specifically, we use $\text{acc}_{\text{benign}}$ to denote the accuracy of the best global model in the benign setting, and $\text{acc}_{\text{drop}}$ represent the accuracy degradation caused by the attack. The attack impact is then defined as $\phi=\text{acc}_{\text{drop}}/\text{acc}_{\text{benign}}\times 100\%$. A higher $\phi$ indicates a more successful attack.
Additionally, we define $\varrho=(|\mathbb{E}^{\text{in}}|-|\mathbb{E}^{\text{in}}\cap\widetilde{\mathbb{E}}^{\text{in}}|)/|\mathbb{E}^{\text{in}}|$ as the percentage of edges that cross the selection boundary caused by the attack, where $\mathbb{E}^{\text{in}}=\{e_{j}\mid w_{j}\geq\gamma\}$ be the top $k\%$ edges selected by the subnetwork before the attack and $\widetilde{\mathbb{E}}^{\text{in}}=\{e_{j}\mid\widetilde{w}_{j}\geq\widetilde{\gamma}\}$ be the top $k\%$ edges selected by the subnetwork after the attack. We refer to this metric as ‘edge cross rate’ ($\varrho$) hereinafter. A higher $\varrho$ implies a more successful attack.
Poisoning Attacks for Comparison. We compare our VEM with six attacks, including Label-flip attack [6], Noise attack, Grad-ascent attack [7], Min-max attack [11], Min-sum attack [11], Rank-reverse attack [15], and we leave the detailed description for each attack in Appendix B.2. For the Noise attack, Grad-ascent attack, Min-max attack, and Min-sum attack, which are optimized in continuous space, we adapt these attacks to FRL by implementing the attack strategy on the importance scores $\theta^{s}$ to achieve a similar attack effect. Note that this represents the most effective implementation feasible within the scope of our study.
@@ 5.1 Comparison with the State-of-the-art Attacks
VEM outperforms existing attacks. Table II shows that our attack outperforms other attacks for all the combinations of the model, AGR, and dataset, which shows an overall 53.23% attack impact and are 3.7×3.7$\times$3.7 × more impactful than others. Specifically, under FRL defense, our attack achieves 61.25% attack impact, which is 3.8×3.8$\times$3.8 × more impactful than others. In the most robust case, Fang integrated FRL, our attack still achieves 41.28% overall attack impact and is 4×4$\times$4 × more impactful than other attacks. In the easiest task, the MNIST dataset, our attack achieves an overall 48% attack impact, which is 10.5×10.5$\times$10.5 × more impactful than others. Experiment results for the CIFAR100 dataset and VGG architecture can be seen in Appendix 14.
@@
5.2 Effect of the Vulnerable Edge Identification and the Optimization
@@
s.t. ${W}_{m}=A\sum_{u=1}^{m}R_{u}^{2d},$
$\widetilde{{W}}_{m}=A\sum_{u=1}^{m}(R_{u}^{2d}P_{u}).$
@@ 5.3 Impact of FRL Parameters
Impact of the percentage of malicious clients. Fig. 10 shows the attack impacts of different percentages of malicious clients varied from 5% to 25% conducted on the MNIST dataset. We note that our attacks outperform the existing attacks by large margins in 47 out of 50 cases, with the gap ranging from 1% to 48%. For all of the AGRs, with an increasing percentage of malicious clients, the impacts of our attack and the gaps between our attack and existing attacks increase.
@@
Impact of client participation rate. In each training round, the server selects a fraction of the clients to participate in training. Fig. 11 (b) shows the attack impact of our attack when the client participation rate ranges from 0.01 to 1 on the MNIST dataset with the FRL framework. We observe that the participation rate has almost no impact on the effectiveness of our attack.
@@ 6.1 Custom Defense
TABLE IV: Comparison of FRL and IBD integrated FRL under VEM attack.
VEM remains effective against IBD. As demonstrated in Table IV, our VEM attack continues to achieve strong attack impact despite the implementation of IBD defenses. This is because the identified vulnerable edges are inherently inconsistent among benign clients, resulting in minimal variation in the average intersection value $\mu_{u}$ for $u\in U$. Additionally, by simply adjusting $\zeta$ in Eq. 7, we can render the malicious updates indistinguishable from benign updates. Consequently, when the IBD defense is applied, not only does it fail to filter out the malicious updates effectively, but it also inadvertently excludes a significant number of benign clients. This will amplify the overall impact of the VEM attack. We also conduct experiments to measure the false positive rate (FPR) and true positive rate (TPR) of the IBD defense. For more detailed experimental results, please refer to Appendix 16.

### v2

@@ 1 Introduction
We evaluate our attack on eight benchmark datasets and integrate eight Byzantine-robust aggregation rules (AGRs) within the FRL framework to test the attack impact of our VEM. The results demonstrate that our attack achieves a significantly greater attack impact compared to state-of-the-art attacks across all combinations of datasets, model architectures, and AGRs, with an overall attack impact of 53.23%, making it 3.7 times more effective than other attacks. Additionally, we developed a custom defense tailored specifically to counter our attack; however, our VEM attack remains effective against this targeted defense. Our key contributions can be summarized as follows:
  • We conduct the first systematic analysis of FRL’s robustness, uncovering a critical vulnerability within the framework. Our findings reveal that FRL does not guarantee the robustness of the system, and there exists some edges that are particularly vulnerable to poisoning attacks. Additionally, we provide a necessary condition to launch a successful attack in the FRL framework.
@@
  • We release the source code at https://github.com/gongzir1/VEM to facilitate future studies in this area.
@@ 2.3 Federated Rank Learning (FRL)
Lottery Tickets Hypothesis and Supermasks. Modern neural networks contain an enormous number of parameters, often leading to overparameterization [33, 34]. To solve this issue, Frankle and Carbin [16] offer an intriguing hypothesis i.e., Lottery Tickets Hypothesis (LTH), which states that within a randomly initialized, dense neural network (supernetwork), there exists a sparse subnetwork such that when trained in isolation, it can achieve the test accuracy as the fully trained original supernetwork within the same number of training iterations. Follow-up work by Ramanujan et al. [35] proposes the edge-popup (EP) algorithm [35] to effectively find subnetworks. Specifically, the EP algorithm assigns different scores to each edge in the supernetwork and uses these scores to determine which edges to retain. The process involves both a forward and a backward pass. During the forward pass, the EP selects the top kof edges with the highest scores. In the backward pass, it updates these scores using the straight-through gradient estimator [36], enabling gradients to pass through non-differentiable operations. The process is conducted iteratively to update the scores and retain the edges that effectively reduce the training loss. By doing so, it identifies a subnetwork that performs as well as the fully trained supernetwork.
@@ 3.2 Motivation
However, we observe that FRL is not inherently robust, and there are specific edges that are particularly vulnerable to poisoning attacks. To verify this argument, we record the reputation difference between each edge and the selection boundary $\gamma$, which is calculated as $\Delta_{j}=|w_{j}-\gamma|$, and $w_{j}$ is the aggregated reputation for edge $e_{j}$ (for $j\in[1,n]$). This metric indicates the minimal modification needed for edge $e_{j}$ to cross the selection boundary, thereby reflecting its vulnerability. A lower value of $\Delta_{j}$ signifies a higher vulnerability. As illustrated in Fig. 3, there is a significant variation in $\Delta_{j}$ (for $j\in[1,n]$), ranging from 0 to 5616 for the MNIST dataset and from 0 to 17208 for the CIFAR10 dataset. (Details of the experiment setting can be seen in Appendix B.1.) Additionally, we use the red dash line to indicate the max adversary damage (e.g., 20% malicious rate). Edges to the left of this line are considered vulnerable, as the required reputation change is smaller than the maximum adversary damage.
@@ 3.3 VEM Overview
Our VEM aims to manipulate malicious updates such that the aggregated global model deviates significantly from their original ones at each training round, thus negatively affecting the global model performance. To effectively achieve this goal and make the FRL framework tackleable, we first provide a formalization of FRL in Section 3.4. Specifically, we expand the 1D ranking into a 2D matrix to reduce the non-differentiable operations involved in the original algorithm.
@@ 3.5 Vulnerable Edges Identification
In this section, we provide a theoretical analysis of the range of vulnerable edges. Intuitively, if the reputation difference $\Delta_{j}$ between edge $e_{j}$ and the selection boundary is smaller than the maximum damage the adversary can cause in one round, we call the edge $e_{j}$ a vulnerable edge. Formally, we have the following theorem.
@@
Implications. Given that the user input is discrete integers ranging from 1 to $n$, the one-round adversary’s damage is limited by a maximum budget i.e., $m(a_{n}-a_{1})$ for $m$ malicious users. Therefore to launch a successful attack for each round (i.e., move edges across the selection boundary), a necessary condition is that the needed reputation change, i.e., $\overline{w}_{\text{min}}-\overline{w}_{v}$ or $\overline{w}_{\text{max}}-\overline{w}_{v}$ should not exceed the one-round adversary’s damage budget. Therefore, the edges that the adversary can alter (i.e., vulnerable edges) are bounded by Eq. (5). Moreover, as $m$ increases, the range of vulnerable edges expands. As shown in Fig. 6, the x-axis denotes the malicious rate and the y-axis indicates the percentage of vulnerable edges calculated based on Theorem 1. As the malicious rate rises, so does the percentage of vulnerable edges. When the malicious rate reaches 0.4, all edges can be considered vulnerable. Even with an extreme scenario where just 1 malicious client (4% malicious rate), there are still 11% and 10% vulnerable edges on the MNIST and CIFAR10 datasets, respectively.
@@
Use Case for Two Estimation Methods. In the IID case, both methods can achieve the desired estimation performance. In the non-IID case, the historical estimation method offers more reliable performance, as it leverages the global ranking from the previous round for estimation. We defer detailed discussion of these two estimation methods in Section 5.3.
@@ 3.6 Optimization
After identifying the vulnerable edges, we aim to target those edges and manipulate their ranking, thus effectively affecting the global model performance. Therefore, we have the following optimization function such that the global models reputation of those vulnerable edges deviate significantly from their original values for each training round. Formally, we have
@@ 4.1 Experimental setup
Datasets. For our main experiment, we use four benchmark datasets widely used in prior works to evaluate our attacks, including MNIST [41], CIFAR10 [42], FashionMNIST [43], and EMNIST [44]. Details are deferred to Appendix B.2. We use an independently and identically distributed dataset (IID) because poisoning FL with IID data is the hardest [10, 11]. We also conduct experiments on CIFAR100 [42], Location30 [45], Purchase100 [46] and Texas100 [46], with the results presented in Appendix C.
@@
Evaluation Metric. We follow previous works [10, 11] and conduct our evaluations by simulating the attack on our local device. We follow [17] to use the attack impact to measure the attack effect. Specifically, we use $\text{acc}_{\text{benign}}$ to denote the accuracy of the best global model in the benign setting, and $\text{acc}_{\text{drop}}$ represent the accuracy degradation caused by the attack. The attack impact is then defined as $\phi=\text{acc}_{\text{drop}}/\text{acc}_{\text{benign}}\times 100\%$. A higher $\phi$ indicates a more successful attack.
Additionally, we define $\rho=(|\mathbb{E}^{\text{in}}|-|\mathbb{E}^{\text{in}}\cap\widetilde{\mathbb{E}}^{\text{in}}|)/|\mathbb{E}^{\text{in}}|$ as the percentage of edges that cross the selection boundary caused by the attack, where $\mathbb{E}^{\text{in}}=\{e_{j}\mid w_{j}\geq\gamma\}$ be the top $k\%$ edges selected by the subnetwork before the attack and $\widetilde{\mathbb{E}}^{\text{in}}=\{e_{j}\mid\widetilde{w}_{j}\geq\widetilde{\gamma}\}$ be the top $k\%$ edges selected by the subnetwork after the attack. We refer to this metric as ‘edge cross rate’ ($\rho$) hereinafter. A higher $\rho$ implies a more successful attack.
Poisoning Attacks for Comparison. We compare our VEM with six attacks, including Label-flip attack [6], Noise attack, Grad-ascent attack [7], Min-max attack [11], Min-sum attack [11], Rank-reverse attack [15], and we leave the detailed description for each attack in Appendix B.3. For Min-max, Min-sum and Fang attacks, we relax the adversary assumption and assume the adversary has full knowledge of benign updates (as required by the attacks). For the Noise attack, Grad-ascent attack, Min-max attack, and Min-sum attack, which are optimized in continuous space, we adapt these attacks to FRL by implementing the attack strategy on the importance scores $\theta^{s}$ to achieve a similar attack effect. Note that this represents the most effective implementation feasible within the scope of our study.
@@ 5.1 Comparison with the State-of-the-art Attacks
VEM outperforms existing attacks. Table II shows that our attack outperforms other attacks for all the combinations of the model, AGR, and dataset, which shows an overall 53.23% attack impact and are 3.7×3.7$\times$3.7 × more impactful than others. Specifically, under FRL defense, our attack achieves 61.25% attack impact, which is 3.8×3.8$\times$3.8 × more impactful than others. In the most robust case, Fang integrated FRL, our attack still achieves 41.28% overall attack impact and is 4×4$\times$4 × more impactful than other attacks. In the easiest task, the MNIST dataset, our attack achieves an overall 48% attack impact, which is 10.5×10.5$\times$10.5 × more impactful than others. Experiment results for CIFAR100, Location30, Purchase100, and Texas100 can be seen in Appendix C.
@@
5.2 Effect of Vulnerable Edge Identification and the Optimization
s.t. ${W}_{m}=A\sum_{u=1}^{m}R_{u}^{2d},$
$\widetilde{{W}}_{m}=A\sum_{u=1}^{m}(R_{u}^{2d}P_{u}).$
@@ 5.2 Effect of Vulnerable Edge Identification and the Optimization
5.3 Evaluation of Estimation Methods
We evaluate the performance of the two estimation methods based on their ability to accurately identify the true vulnerable edges. Specifically, we first calculate the estimated $\overline{W}$ and get the vulnerable edges set using these two methods separately $\mathbb{V}_{\text{alt / hist}}$. We then compare these sets with the true set of vulnerable edges $\mathbb{V}_{\text{true}}$ which is determined using full knowledge (ideal case). Then we calculate the estimation accuracy (i.e., $|\mathbb{V}_{\text{alt / hist}}\cap\mathbb{V}_{\text{true}}|/|\mathbb{V}_{\text{true}}|$) to measure the estimation performance. A higher value indicates a more accurate estimation. As illustrated in Fig. 8, where the y-axis represents the estimation accuracy and the x-axis depicts training iteration. In IID case, both methods achieve a estimation accuracy above 98%, and in non-IID case, the historical estimation’s accuracy is above 95%.
Figure 8: Estimation accuracy during the FRL training.
This high estimation accuracy for the historical method in non-IID case can be attributed primarily to two key factors: Firstly, as shown in Theorem 1, we only need to estimate the aggregated reputation of benign clients to identify vulnerable edges. This aggregated result makes our method less sensitive to individual client updates. Secondly, although the historical estimation method relies on the assumption that the aggregated client reputation remains relatively stable across rounds. This assumption is reasonable because, in each training round, clients receive the global ranking and train for only a few epochs (e.g., 2). As a result, $R_{u}(t)$ remains relatively close to $R_{g}(t-1)$. Significant deviations would also negatively impact global model convergence. Additionally, as the model converges, this approximation becomes more accurate even in non-IID settings.
@@ 5.4 Impact of FRL Parameters
Here, we provide a detailed explanation of why our attack remains effective in the non-IID setting. As discussed in Section 5.3, our historical estimation achieves the desired performance, ensuring that the non-IID nature of the data does not significantly hinder our attack. Moreover, our attack performs even better as the non-IID rate increases. This is because higher non-IID levels weaken the consensus among benign clients, leading to a decrease in $\overline{w}_{\text{max}}$ and an increase in $\overline{w}_{\text{min}}$ in Eq. 5. This expansion of the vulnerable edge range reduces our attack’s dependence on specific edges, as we can effectively identify most of them. Additionally, diminished consensus among benign clients makes edges more susceptible to manipulation, further increasing the success rate of our attack.
Impact of the percentage of malicious clients. Fig. 10 shows the attack impacts of different percentages of malicious clients varied from 5% to 25% conducted on the MNIST dataset. We note that our attacks outperform the existing attacks by large margins in 47 out of 50 cases, with the gap ranging from 1% to 48%. For all of the AGRs, with an increasing percentage of malicious clients, the impacts of our attack and the gaps between our attack and existing attacks increase. Table. IV shows attack impact on more datasets with malicious rate ranking from 10% to 25%.
CIFAR10 FashionMNIST Location30 Purchase100 Texas100
(Conv8) (LeNet) (FC) (FC) (FC)
10% 64.51 39.10 88.01 95.65 51.56
15% 66.41 42.47 89.20 98.43 52.53
20% 71.73 50.77 90.02 99.80 67.19
25% 81.69 75.84 93.22 99.80 88.21
TABLE IV: Attack impact (%) of VEM on more datasets and model architectures with different malicious rates under FRL framework.
@@
Impact of client participation rate. In each training round, the server selects a fraction of the clients to participate in training. Fig. 11 (b) shows the attack impact of our attack when the client participation rate ranges from 0.01 to 1 on the MNIST dataset with the FRL framework. We observe that the participation rate has almost no impact on the effectiveness of our attack, which demonstrates the scalability of our attack.
@@ 6.1 Custom Defense
TABLE V: Comparison of FRL and IBD integrated FRL under VEM attack.
VEM remains effective against IBD. As demonstrated in Table V, our VEM attack continues to achieve strong attack impact despite the implementation of IBD defenses. This is because the identified vulnerable edges are inherently inconsistent among benign clients, resulting in minimal variation in the average intersection value $\mu_{u}$ for $u\in U$. Additionally, by simply adjusting $\zeta$ in Eq. 7, we can render the malicious updates indistinguishable from benign updates. Consequently, when the IBD defense is applied, not only does it fail to filter out the malicious updates effectively, but it also inadvertently excludes a significant number of benign clients. This will amplify the overall impact of the VEM attack. We also evaluate the False Positive Rate (FPR) and True Positive Rate (TPR) of the IBD defense under different values of hyper-parameter $\zeta$ in Eq. (7). As shown in Fig. 14, as the value of $\zeta$ increases, the FPR decreases while the TPR increases. However, it still has a relatively high FPR, e.g., more than 20% for the Fashion-MNIST dataset and around 20% for the CIFAR10 dataset, indicating that some benign updates are filtered out by the defense. Therefore, the IBD-integrated FRL is not yet sufficiently effective against our VEM attack.